OpenStudy (anonymous):

how do u find trig values by using theta + 2kPi= theta ie of 1390

OpenStudy (anonymous):

Always draw the trig circle. Obviously the trig value of an angle theta is the same as the trig value of the same angle theta + 2pi because both angles lie in the same point of the circle. This holds true for any number of 2pi rotations you take. Is this what you were looking for?

OpenStudy (anonymous):

how do i find it without drawing a circle?

OpenStudy (anonymous):

Its simple logic. 2kpi means 360 times K. So if youre facing some way, and you take any number of 360% spins, you will face the same way in the end!

OpenStudy (anonymous):

Sorry, 360 degrees.

OpenStudy (phi):

if you divide by 2pi and look at the remainder, that is the equivalent angle type 1390/(2pi)= in the google search window that tells you how many times 2pi goes into 1390. subtract off the integer (the whole number) to get the remainder

OpenStudy (anonymous):

i dont know how to break the angle into \[\theta +2k\]

OpenStudy (phi):

btw, if 1390 is in degrees, you divide by 360, and take the remainder

OpenStudy (anonymous):

Just subtract 2*pi from it as many times as you want, and stop if youre about to go negative.

OpenStudy (anonymous):

this looks ultra confusing because you have a \(\pi\) on the left and what looks like degrees on the right. you have to use either degrees or radians, cannot mix them

OpenStudy (anonymous):

i get 30 plus 2k pi what do i insert in k?

OpenStudy (phi):

Here is an example. say you start with 3pi radians (we know the answer should be pi) (3pi/2pi)= 1.5 subtract the 1 and you get 0.5 as the remainder (that means 1/2 the way around the circle). multiply by 2pi to get the final answer: 2*0.5pi= pi so (pi + 2pi) = 3pi (here k=1)

OpenStudy (anonymous):

ok thank u all!

OpenStudy (phi):

you should get (1390 rads)/(2pi)= 221.225371 so theta= 2pi*0.225371 + 221*2pi or 0.45pi + 221*2pi, k= 221

OpenStudy (phi):

*0.45 is rounded